# One-to-N Backdoor Attack in 3D Point Cloud via Spherical Trigger

### v1

@@ 4.2.3 Performance on Benchmark Datasets
The experimental results demonstrate the high effectiveness of our one-to-N backdoor attack across all evaluated configurations. The dual-sphere trigger design consistently outperforms the single-sphere variant, achieving superior attack success rates as detailed in previous sections. More importantly, we observe a remarkably consistent performance pattern across the four target classes (T1-T4) within each experimental setting. The ASR values for different targets show minimal variation, indicating that our spatial configuration framework establishes equally reliable backdoor associations for all targets without exhibiting bias towards any specific class. This uniform performance across multiple targets, combined with the consistently high success rates (predominantly exceeding 90% and often approaching 100%), validates the robustness and scalability of our approach in creating separable and effective trigger-to-target mappings.
Correspondingly, the Clean Accuracy (ACC) analysis reveals that the backdoor implantation has minimal impact on the model’s primary functionality. The ACC values remain stable across different trigger designs for each model-dataset combination, demonstrating that the embedded backdoors do not compromise the model’s utility on clean samples. The models’ maintained accuracy on clean inputs, coupled with their high vulnerability to triggered samples, highlights a clear performance disparity and underscores the stealthiness of our attack—the backdoored models preserve their expected performance on benign inputs while being susceptible to targeted manipulation when exposed to the trigger patterns.
@@ 4.2.4 Effectiveness with Different $N$ Values
To comprehensively evaluate the scalability of the one-to-N attack methodology, we conduct experiments on the PointNet model with varying numbers of target classes ($N=1,2,3,4$) while maintaining a consistent poisoning ratio of 0.01 per target across all configurations. This systematic investigation, summarized in Table 3, aims to understand how the attack performance scales with increasing complexity of multi-target backdoor implantation while examining the relationship between the total poisoning ratio, attack success rate, and model utility.
@@ 4.2.5 Resistance to SOR
Table 4 evaluates the effectiveness of Statistical Outlier Removal (SOR) defense against our proposed backdoor attack mechanism. Our approach employs the insertion of additional points in the form of small spheres around the point cloud periphery, which makes SOR the most relevant and effective countermeasure since it specifically targets statistical outliers. To rigorously test SOR’s defensive capability under the most challenging conditions, we selected the 1% poisoning ratio per target on the ShapeNetPart dataset. This configuration represents an extreme attack scenario where the backdoor attack achieves nearly 100% Attack Success Rate (ASR) in the absence of any defense mechanisms. By testing SOR defense under these optimal attack conditions, we can more clearly demonstrate its effectiveness and limitations.
@@ 5 Conclusion
This work answers the critical question of how to realize one-to-N backdoor attacks in 3D point clouds by presenting a novel framework centered on configurable spherical triggers. The framework’s feasibility is theoretically established in the Appendix, where our analysis proves that the poisoned model can associate distinct spherical trigger configurations with different target labels based on their spatial properties. This theoretical conclusion is further validated through rigorous experimentation, demonstrating that multi-target backdoor attacks are not only feasible but pose a severe and flexible threat.
The primary contributions of this work are twofold: first, it establishes the first comprehensive framework for one-to-N backdoor attacks in 3D point clouds, instantiated through two concrete implementations based on configurable spherical triggers; second, it provides extensive experimental validation and theoretical proofs that rigorously confirm the effectiveness of the spherical trigger approaches. These contributions collectively open a new research frontier by proving the viability of parameter configuration as a discriminative mechanism in 3D backdoor attacks, paving the way for even more flexible, stealthy, and defense-resistant one-to-N attacks in 3D domains.
@@ A.2 Key Assumptions and Simplifications
(2)Uniform Class Distribution: We assume approximately uniform class distribution in benign training data, enabling the simplification $\sum_{j=1}^{H-M}K(\bm{X},\bm{X}_{j})\cdot\mathbb{I}(y_{j}=t_{i})\approx\sum_{j=1}^{(H-M)/C}K(\bm{X},\bm{X}_{j})$.
@@ A.3.2 Proof of Lemma 2
Let $C=C(\bm{X}^{\prime}_{R_{0}})$, $A=A(\bm{X}^{\prime}_{R_{0}})$, $B_{0}=B(\bm{X}^{\prime}_{R_{0}})$, and $d=\|R-R_{0}\|$. The main trend of $\varphi_{t_{i}}(\bm{X}^{\prime}_{R})$ is captured by the function:

### v2

@@ 4.2.3 Performance on Benchmark Datasets
The experimental results demonstrate the high effectiveness of the STONE framework attack across all evaluated configurations. The dual-sphere trigger design consistently outperforms the single-sphere variant, achieving superior attack success rates as detailed in previous sections. More importantly, we observe a remarkably consistent performance pattern across the four target classes (T1-T4) within each experimental setting. The ASR values for different targets show minimal variation, indicating that our spatial configuration framework establishes equally reliable backdoor associations for all targets without exhibiting bias towards any specific class. This uniform performance across multiple targets, combined with the consistently high success rates (predominantly exceeding 90% and often approaching 100%), validates the robustness and scalability of our approach in creating separable and effective trigger-to-target mappings.
Correspondingly, the Clean Accuracy (ACC) analysis reveals that the backdoor implantation has minimal impact on the model’s primary functionality. The ACC values remain stable across different trigger designs for each model-dataset combination, demonstrating that the embedded backdoors do not compromise the model’s utility on clean samples. The models’ maintained accuracy on clean inputs, coupled with their high vulnerability to triggered samples, highlights a clear performance disparity and underscores the stealthiness of STONE—the backdoored models preserve their expected performance on benign inputs while being susceptible to targeted manipulation when exposed to the trigger patterns.
@@ 4.2.4 Effectiveness with Different $N$ Values
To comprehensively evaluate the scalability of the STONE framework, we conduct experiments on the PointNet model with varying numbers of target classes ($N=1,2,3,4$) while maintaining a consistent poisoning ratio of 0.01 per target across all configurations. This systematic investigation, summarized in Table 3, aims to understand how the STONE framework’s performance scales with increasing complexity of multi-target backdoor implantation while examining the relationship between the total poisoning ratio, attack success rate, and model utility.
@@ 4.2.5 Resistance to SOR
Table 4 evaluates the effectiveness of Statistical Outlier Removal (SOR) defense against the STONE framework. The STONE framework employs the insertion of additional points in the form of small spheres around the point cloud periphery, which makes SOR the most relevant and effective countermeasure since it specifically targets statistical outliers. To rigorously test SOR’s defensive capability under the most challenging conditions, we selected the 1% poisoning ratio per target on the ShapeNetPart dataset. This configuration represents an extreme attack scenario where the backdoor attack achieves nearly 100% Attack Success Rate (ASR) in the absence of any defense mechanisms. By testing SOR defense under these optimal attack conditions, we can more clearly demonstrate its effectiveness and limitations.
@@ 5 Conclusion
This work establishes STONE, the first comprehensive framework that demonstrates how to realize one-to-N backdoor attacks in 3D point clouds using configurable spherical triggers. The framework’s feasibility is theoretically grounded in Neural Tangent Kernel analysis, which shows that poisoned models can learn to associate distinct spatial trigger configurations with different target labels. This theoretical foundation is extensively validated through rigorous experiments across multiple datasets and architectures, confirming that multi-target backdoor attacks represent a severe and adaptable threat to 3D vision systems.
Our work provides both a theoretical basis and a practical implementation for this new attack paradigm, establishing spherical triggers as an effective mechanism for dynamic target control in 3D domains. The STONE framework opens a new research direction by proving the viability of spatial configuration as a discriminative mechanism, paving the way for developing more flexible, stealthy, and defense-resistant one-to-N attacks.
@@ A.2 Key Assumptions and Simplifications
(2)Uniform Class Distribution: We assume approximately uniform class distribution in benign training data, enabling the simplification $\sum_{j=1}^{H-M}K(\bm{X},\bm{X}j)\cdot\mathbb{I}(y_{j}=t_{i})\approx\sum{j=1}^{(H-M)/C}K(\bm{X},\bm{X}_{j})$, where $\mathbb{I}(y_{j}=t_{i})$ is an indicator function that equals 1 if sample $j$ belongs to class $t_{i}$ and 0 otherwise.
@@ A.3.2 Proof of Lemma 2
where $C=C(\bm{X}^{\prime}_{R_{0}})$, $A=A(\bm{X}^{\prime}_{R_{0}})$, $B_{0}=B(\bm{X}^{\prime}_{R_{0}})$, and $d=\|R-R_{0}\|$. The main trend of $\varphi_{t_{i}}(\bm{X}^{\prime}_{R})$ is captured by the function: